OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: 07_08_LQ_4.gif Which equation gives the cost, y, in terms of the number of pages, x?

OpenStudy (anonymous):

The Table is Number of pages: 50|100|150|200 Cost : 1.25|1.85|2.45|3.05|

OpenStudy (anonymous):

A. y = 83.3x + 0.65 B. y = 0.012x + 1.25 C. y = 0.012x + 49.985 D. y = 0.012x + 0.65

OpenStudy (smileyxl3):

0.012x + .65